2.2. Complete the table below for the Third National Bank. You have to distinguish between a bank's assets and bank's liabilities.

The figures in the table below are for the Third National Bank. All figures are in thousands of dollars.

Assets

Liabilities and Net Worth

Stock Shares

$    420  

$

_____

$

_____

Reserves

25

_____

_____

Property

300

____

_____

Securities

100

____

_____

Loans

100

____

_____

Demand Deposits

105

____

_____

2.3. What is the total assets of this bank? Explain the basics of this bank’s balance sheet.

2.4. If the required reserve ratio for the Third National Bank is 10 percent, what is the monetary multiplier?

Recall, to calculate you have to use the formula: Monetary Multiplier = 1÷Required Reserve Ratio. The money multiplier is a key measure in banking that helps to predict the money supply that will be available to drive economic growth. As you can see from the formula, if the reserve requirement is 20%, the money multiplier will be 1 divided by 0.2, which is 5. We can then use the money multiplier multiplied by the excess reserves to determine the maximum checkable-deposit creation that will be provided by the new money entering the system.

2.5. If the monetary multiplier is 4, what is the required reserve ratio? Describe how and identify by what amount the Third National Bank can create money in the economy.

Recall that, generally, bank creates money in a typical economy by making loans. The Fed sets the reserve requirement (the required reserve ratio) that directly affects the amount of money creation.
Answer #1

2.3 Total assets are $525

Assets Liabilities and Net Worth
Stock Shares 420
Reserves 25
Property 300
Securities 100
Loans 100
Demand Deposits 105
Total 525 525

2.4 Money multiplier = 1 / required reserve ratio = 1 / 0.10 = 10

2.5 If money multiplier = 4, then required reserve ratio = 1 / 4 = 0.25